# LaCATODA 2026: The 10th Linguistic and Cognitive Approaches to Dialog Agents Workshop at the 40th AAAI conference, Janurary 20–27, 2026, Singapore Expo

### v3

@@ Abstract
We present CID-GraphRAG (Conversational Intent-Driven Graph Retrieval Augmented Generation), a novel framework that addresses the limitations of existing dialogue systems in maintaining both contextual coherence and goal-oriented progression in multi-turn customer service conversations. Unlike traditional RAG systems that rely solely on semantic similarity (Conversation RAG) or standard knowledge graphs (GraphRAG), CID-GraphRAG constructs dynamic intent transition graphs from goal-achieved historical dialogues and implements a dual-retrieval mechanism that adaptively balances intent-based graph traversal with semantic search. This approach enables the system to simultaneously leverage both conversational intent flow patterns and contextual semantics, significantly improving retrieval quality and response quality. In extensive experiments on real-world customer service dialogues, we demonstrated that CID-GraphRAG significantly outperforms both semantic-based and intent-based baselines across automatic metrics, LLM-as-Judge evaluations and human evaluations, with relative gains of $11.4\%$ in BLEU, $4.9\%$ in ROUGE, $5.9\%$ in METEOR, and most notably, a $58.0\%$ improvement in response quality according to LLM-as-judge evaluations. These results demonstrate that integrating intent transition structures with semantic retrieval creates a synergistic effect that neither approach achieves independently, establishing CID-GraphRAG as an effective framework for addressing the challenges of maintaining contextual coherence and goal-oriented progression in knowledge-intensive multi-turn dialogues.
@@ 1 Introduction
Large Language Models (LLMs) struggle with domain-specific multi-turn dialogues despite Retrieval-Augmented Generation (RAG) enhancements [Lewis2020]. Customer service dialogues present unique challenges: temporal dependencies creating complex reference chains across turns, and dynamically evolving conversation objectives that traditional RAG approaches fail to address effectively.
Current research follows two separate trajectories: Conversation RAG methodologies [researchgate2025learning, ye2024boosting] optimize dialogue history representation through semantic matching, while GraphRAG approaches [graphrag2025, graphrag2024, zhang2025credible] leverage knowledge graphs for reasoning over static domain knowledge. Both have fundamental limitations—Conversation RAG lacks structural guidance for conversation flow, while GraphRAG fails to model dynamic intent transitions essential for coherent multi-turn dialogues.
We propose CID-GraphRAG (Conversational Intent-Driven Graph Retrieval Augmented Generation), a novel framework that integrates intent transition graphs with semantic similarity retrieval. Our approach leverages goal-achieved historical conversations to model conversation dynamics through both semantic content and conversation flow, enabling more coherent and goal-oriented responses. The key contributions of our work include:
  • Intent transition graph construction: automatically building dynamic intent transition graphs from goal-achieved historical dialogues, systematically cataloging hierarchical intents from both system and user utterances to create a network of effective conversational trajectories.
  • Dual-pathway adaptive retrieval: balancing intent transition graph traversal with semantic similarity search, through an adaptive weighting mechanism. This integration enables the system to simultaneously leverage conversation structure and contextual semantics, addressing the limitations of single-signal approaches in complex multi-turn dialogues.
@@ 2 Related Works
RAG enhances Large Language Models (LLMs) by integrating external knowledge sources, improving response accuracy and informativeness in dialogue systems [Lewis2020]. Recent advancements include memory-based approaches like DH-RAG [zhang2025dhrag] and SELF-multi-RAG [self-rag], evaluation frameworks like MTRAG [katsis2025mtrag], and specialized applications combining knowledge graphs for customer service [xu2024retrieval] and real-time problem detection with FAQ retrieval [agrawal2024beyond].
GraphRAG enhances traditional RAG by integrating knowledge graphs for structured semantic retrieval and multi-hop reasoning [graphrag2025], better capturing contextual dependencies in multi-turn dialogues [graphrag2024]. Key advantages include enhanced retrieval accuracy through preserved structural relationships [xu2024retrieval], effective processing of multi-hop queries requiring cross-fragment reasoning [graphrag2025, zhang2025credible], context continuity maintenance via entity relationship tracking [xu2024retrieval], and improved answer interpretability through visualized reasoning paths [linders2025knowledge].
Conversation RAG addresses multi-turn dialogue challenges in customer service, where user queries often depend on historical interactions [researchgate2025learning, ye2024boosting].Key advancements include context-aware frameworks like DH-RAG [zhang2025dhrag] and SELF-multi-RAG [self-rag]; structured retrieval approaches such as KG-RAG [xu2024retrieval]; query decomposition techniques like Collab-RA [collab-rag]; and specialized evaluation benchmarks including RAD-Bench [kuo2024radbench], MTRAG [katsis2025mtrag], and ChatQA [chatqa]. While these approaches have advanced independently, no unified solution simultaneously addresses both contextual coherence and structured reasoning challenges.
@@ 3 Methodology
CID-GraphRAG innovatively integrates Conversation RAG and GraphRAG to address the challenge of simultaneously maintaining contextual coherence and goal-oriented progression in customer service dialogues. Our methodology consists of three key components: intent graph construction, adaptive dual-pathway retrieval, and structured response generation.
@@ 3.1 System Overview
Figure 1: The detailed framework of CID-GraphRAG. The CID-GraphRAG consists of two phases: (1) a construction phase that builds an intent graph from goal-achieved conversations, and (2) an inference phase that identifies user and assistant intents from current dialogue turn, retrieves high-quality examples from both intent-based and semantic-based pathway via adaptive weighting mechanism, and uses LLM for structured response generation.
@@ 3.2.1 Intent Recognition
Our intent recognition process employs a two-stage classification approach powered by LLM, first identifying primary intents, then identifying secondary intents within that category (illustrated in Appendix A). This can improve intent classification accuracy compared to direct fine-grained classification.
@@ 3.2.2 Graph Construction
After intent recognition, we organize intents into a graph structure that captures conversation dynamics, as illustrated in Figure 2). Our intent graph contains three node types:
@@
These nodes are connected through hierarchical relations (has_child_intent), pairing relations (in_pair_with), transition relations (next_intent), and dialogue anchoring (has_conversation). This graph structure enables the system to navigate conversation trajectories that have historically led to successful outcomes.
Figure 2: Internal structure of the CID-Graph.The graph comprises distinct primary intent and secondary intent nodes and conversation nodes. Key relations include hierarchical, pairing, transition, and dialogue anchoring.
@@ 3.3 Inference Phase
During the inference phase, we propose a dual-layer retrieval mechanism that integrates intent transition retrieval and semantic similarity retrieval. This dual-layer approach enables mutually-assisted adaptive retrieval, enhancing both retrieval effectiveness and efficiency.
@@ 3.3.1 Dual-Pathway Retrieval
The core innovation of CID-GraphRAG is its dual-pathway retrieval mechanism that dynamically balances intent structure with semantic context. By simultaneously leveraging conversation flow patterns from successful historical dialogues and contextual semantics from the current interaction, this approach enables mutually-assisted adaptive retrieval that enhances both retrieval relevance and goal orientation.
@@
    Graph retrieval: retrieving candidate intents based on the normalized co-occurrence frequencies:
@@
Semantic-Based Retrieval Path: We integrate Conversation RAG semantic retrieval within the intent graph (details in Appendix C):
@@ 3.3.2 Adaptive Integration
These top-k dialogues and their intents serve as few-shot examples for response generation.
@@ 4.1.1 Dataset
We conducted experiments on a real-world customer service dataset consisting of 268 conversations between representatives and drivers regarding vehicle sticker issues. The dataset exhibits strong task-orientedness and multi-turn complexity, covering multiple subtopics including compliance consultation, violation inquiries, appointment scheduling, service coordination, etc. All dialogues were manually selected and annotated by domain experts with extensive experience in customer service. A dataset example is shown in Appendix D.
The dataset contains 1,574 total dialogue turns with an average of 6.1 turns per dialogue (ranging from 2 to 24 turns) and 21.98 words per utterance on average. All dialogues have been rigorously anonymized to protect user privacy. Following standard practice, we partitioned the dataset into training set ($80\%$, 214 dialogues with 1,299 conversation turns), validation set ($10\%$, 27 dialogues with 149 turns), and test set($10\%$, 27 dialogues with 126 turns).
@@ 4.1.2 Evaluation Metrics
Automatic Metrics: We calculated BLEU, ROUGE, METEOR and BERTSCORE between system response and ground truth responses, as well as between retrieved responses and ground truth.
LLM-as-Judge Evaluation: We utilized Claude 3.7 Sonnet [sonnet_aws] as an independent evaluator to assess both retrieved response quality and generated responses. For each evaluation instance, the LLM was provided with the complete dialogue history, the current user query, anonymized system responses generated by different systems, and the ground truth response for reference. For each comparison, LLM compared outputs from all systems across five dimensions (relevance, helpfulness, style consistency, contextual appropriateness, and professionalism) and provided numerical ratings (1-10) for each system, with the highest-rated system designated as "winner" for that example. For retrieved quality evaluation, we excluded the "professionalism" dimension, and removed ground truth responses in the LLM prompt. The complete prompt template for reponse generation is provided in Appendix E.
@@ 4.2 Hyperparameter Analysis
Our empirical analysis reveals two key findings regarding the parameter $\alpha$: (1) the configuration with $\alpha=0.1$ consistently achieves optimal performance across all evaluation metrics, with performance degradation as intent weight increases; and (2) the configurations with $\alpha\geq 0.5$ yield identical retrieval results, suggesting a saturation threshold beyond which intent-based retrieval dominates the ranking process. To validate these automatic evaluation results, we performed LLM-as-Judge evaluation and confirmed the results are aligned with our findings.
@@ 4.3 Evaluation Results
Table 2: Performance comparison across all metrics. Bold indicates best performance for each metric. The percentage values in parentheses show CID-GraphRAG’s relative improvement over the second-best system.
Metric Retrieval Quality Response Generation
@@
Table 2 presents comprehensive results comparing CID-GraphRAG with baselines for retrieval quality and response generation respectively.
Retrieval Quality: CID-GraphRAG demonstrates consistent superiority in retrieval quality across all metrics. For automatic metrics, it achieves relative gains of 2.9% in BLEU-4, 5.2% in ROUGE-L, and a substantial 7.0% in METEOR compared to the second-best system. The LLM-as-judge evaluation reveals even more significant improvements, with CID-GraphRAG outperforming other approaches across all dimensions, particularly in relevance (+12.5%) and helpfulness (+9.3%), demonstrating that our dual-pathway approach retrieves more useful examples.
@@
LLM-as-Judge evaluation of generated responses reveals significant advantages. CID-GraphRAG achieves the highest scores across all dimensions, with notable improvements in style consistency (+4.9%) and professionalism (+4.3%). Most importantly, CID-GraphRAG records the highest number of wins in head-to-head comparisons, representing a 58% improvement over the second best (Conversation RAG), as illustrated in Appendix F.
@@ 4.4 Discussion
Comparing human and LLM-as-Judge evaluations reveals interesting differences in system rankings. While both confirm CID-GraphRAG’s superiority, LLM-as-Judge identified Direct LLM as competitive in style consistency (7.83) and relevance (7.81), whereas human evaluators consistently ranked Intent RAG second, particularly for style consistency (9.38) and professionalism (9.03). This suggests domain experts place higher value on intent-guided responses than LLM evaluators, especially for maintaining professional communication standards in customer service contexts.
While CID-GraphRAG delivers superior retrieval and response quality, it introduces additional computational overhead primarily from LLM-based intent recognition, resulting in higher token consumption and latency compared to baseline approaches. However, our analysis in Appendix G demonstrates that replacing LLM-based intent recognition with a fine-tuned lightweight classification model can reduce latency from approximately 3 seconds to 0.01 seconds, bringing CID-GraphRAG’s total cost and response time to levels comparable with direct LLM generation while preserving most quality advantages.

### v4

# LaCATODA 2026: The 10th Linguistic and Cognitive Approaches to Dialog Agents Workshop at the 40th AAAI conference, January 20–27, 2026, Singapore Expo

@@ Abstract
We present CID-GraphRAG (Conversational Intent-Driven Graph Retrieval-Augmented Generation), a novel framework that addresses the limitations of existing dialogue systems in maintaining both contextual coherence and goal-oriented progression in multi-turn customer service conversations. Unlike traditional RAG systems that rely solely on semantic similarity or static knowledge graphs, CID-GraphRAG constructs intent transition graphs from goal-achieved historical dialogues and implements a dual-retrieval mechanism that balances intent-based graph traversal with semantic search. This approach enables the system to simultaneously leverage both conversational intent flow patterns and contextual semantics, significantly improving retrieval quality and response quality. In extensive experiments on real-world customer service dialogues, we demonstrated that CID-GraphRAG significantly outperforms both semantic-based and intent-based baselines across automatic metrics, LLM-as-a-Judge evaluations and human evaluations, with relative gains of $11.4\%$ in BLEU, $4.9\%$ in ROUGE, and $5.9\%$ in METEOR. Most notably, CID-GraphRAG achieves a $57.9\%$ improvement in response quality according to LLM-as-a-Judge evaluations. These results demonstrate that integrating intent transition structures with semantic retrieval creates a synergistic effect that neither approach achieves independently, establishing CID-GraphRAG as an effective framework for real-world multi-turn dialogue systems in customer service and other knowledge-intensive domains.
@@ 1 Introduction
Large Language Models (LLMs) face significant challenges in domain-specific multi-turn dialogues despite Retrieval-Augmented Generation (RAG) enhancements [1]. Customer service dialogues present unique challenges: temporal dependencies creating complex reference chains across turns, and dynamically evolving conversation objectives that traditional RAG approaches fail to address effectively.
Current research follows two separate trajectories - Conversation RAG methodologies [2, 3] optimize dialogue history representation through semantic matching, while GraphRAG approaches [4, 5, 6] leverage knowledge graphs for reasoning over static domain knowledge. Both have fundamental limitations—Conversation RAG lacks structural guidance for conversation flow, while GraphRAG fails to model intent transitions essential for coherent multi-turn dialogues.
We propose CID-GraphRAG (Conversational Intent-Driven Graph Retrieval-Augmented Generation), a novel framework that integrates intent transition graphs with semantic similarity retrieval. Our approach leverages goal-achieved historical conversations to model conversation dynamics through both semantic content and conversation flow, enabling more coherent and goal-oriented responses. The key contributions of our work include:
  • Intent transition graph construction: automatically building intent transition graphs from goal-achieved historical dialogues, systematically cataloging hierarchical intents from both assistant and user utterances to create a network of effective conversational trajectories.
  • Dual-pathway retrieval: balancing intent transition graph with semantic similarity search, through a weighting mechanism. This integration enables the system to simultaneously leverage conversation structure and contextual semantics, addressing the limitations of single-signal approaches in complex multi-turn dialogues.
@@ 2 Related Works
RAG enhances Large Language Models (LLMs) by integrating external knowledge sources, improving response accuracy and informativeness in dialogue systems [1]. Recent advancements include memory-based approaches like DH-RAG [7] and SELF-multi-RAG [8], evaluation frameworks like MTRAG [9], and specialized applications combining knowledge graphs for customer service [10] and real-time problem detection with FAQ retrieval [11].
GraphRAG enhances traditional RAG by integrating knowledge graphs for structured semantic retrieval and multi-hop reasoning [4], better capturing contextual dependencies in multi-turn dialogues [5]. Key advantages include enhanced retrieval accuracy through preserved structural relationships [10], effective processing of multi-hop queries requiring cross-fragment reasoning [4, 6], context continuity maintenance via entity relationship tracking [10], and improved answer interpretability through visualized reasoning paths [12].
Conversation RAG addresses multi-turn dialogue challenges in customer service, where user queries often depend on historical interactions [2, 3]. Key advancements include context-aware frameworks like DH-RAG [7] and SELF-multi-RAG [8]; structured retrieval approaches such as KG-RAG [10]; query decomposition techniques like Collab-RA [13]; and specialized evaluation benchmarks including RAD-Bench [14], MTRAG [9], and ChatQA [15]. While these approaches have advanced independently, no unified solution simultaneously addresses both contextual coherence and structured reasoning challenges.
Figure 1: The detailed framework of CID-GraphRAG. The CID-GraphRAG consists of two phases: (1) a construction phase that builds an intent graph from goal-achieved conversations, and (2) an inference phase that identifies user and assistant intents from current dialogue turn, retrieves high-quality examples from both intent-based and semantic-based pathway via weighting mechanism, and uses LLM for structured response generation.
@@ 3 Methodology
CID-GraphRAG innovatively integrates Conversation RAG and GraphRAG to address the challenge of simultaneously maintaining contextual coherence and goal-oriented progression in customer service dialogues. Our methodology consists of three key components: intent graph construction, dual-pathway retrieval, and structured response generation.
@@ 3.2.1 Intent Recognition
Our intent recognition process employs a two-stage classification approach powered by LLM, first identifying primary intents, then identifying secondary intents within that category. This hierarchical approach improves intent classification accuracy compared to direct fine-grained classification. Appendix A provides concrete examples demonstrating how this framework disambiguates utterances with identical wording but different contextual meanings.
The complete classification procedure is formalized in Algorithm 1, where we use Claude 3.7 Sonnet to classify both assistant and user utterances in each dialogue turn.
@@ 3.2.2 Graph Construction
After intent recognition, we organize intents into a graph structure that captures conversation dynamics, as illustrated in Figure 2. Our intent graph contains three node types:
@@
Input: Dialogue turn $T$ with current utterances and dialogue history
foreach participant $P\in\{\text{assistant},\text{user}\}$ do
  Stage 1: Primary Intent Classification
 Construct prompt with:
  - Domain context
  - Dialogue history
  - Current dialogue turn
  - List of primary intents
 Send prompt to Claude 3.7 Sonnet
  Extract primary intent $I^{P}_{1}$ from response
 Stage 2: Secondary Intent Classification
 Construct prompt with:
  - Domain context
  - Dialogue history
  - Current dialogue turn
  - Primary intent $I_{1}^{P}$ determined in Stage 1
  - Candidate secondary intents specific to $I_{1}^{P}$
 Send prompt to Claude 3.7 Sonnet
  Extract secondary intent $I^{P}_{2}$ from response
end foreach
Algorithm 1 Two-Stage Intent Classification for Dialogue Turns
Figure 2: Internal structure of the CID-Graph. The graph comprises distinct primary intent nodes, secondary intent nodes and conversation nodes. Key relations include hierarchical, pairing, transition and dialogue anchoring.
These nodes are connected through hierarchical relations (has_child_intenthas\_child\_intent, has_parent_intenthas\_parent\_intent), pairing relations (in_pair_within\_pair\_with), transition relations (next_intentnext\_intent), and dialogue anchoring (has_conversationhas\_conversation). This graph structure enables the system to navigate conversation trajectories that have historically led to successful outcomes.
@@ 3.3 Inference Phase
During the inference phase, we propose a dual-layer retrieval mechanism that integrates intent transition retrieval and semantic similarity retrieval. This dual-layer approach enables mutually-assisted retrieval, enhancing both retrieval effectiveness and efficiency.
@@ 3.3.1 Dual-Pathway Retrieval
The core innovation of CID-GraphRAG is its dual-pathway retrieval mechanism that balances intent structure with semantic context. By simultaneously leveraging conversation flow patterns from successful historical dialogues and contextual semantics from the current interaction, this approach enables mutually-assisted retrieval that enhances both retrieval relevance and goal orientation.
@@
    Graph retrieval: retrieving candidate intents based on the normalized co-occurrence frequencies. To ensure compatibility with semantic similarity scores, we normalize frequencies using:
@@
Semantic-Based Retrieval Path: We integrate semantic retrieval within the intent graph (details in Appendix C):
@@ 3.3.2 Weighted Integration
These top-k dialogues serve as few-shot examples for response generation.
@@ 4.1.1 Dataset
We conducted experiments on a real-world customer service dataset consisting of 268 conversations between representatives and drivers regarding vehicle sticker issues. The dataset exhibits strong task-orientedness and multi-turn complexity, covering multiple subtopics including compliance consultation, violation inquiries, appointment scheduling, service coordination, etc. All dialogues were manually selected and annotated by domain experts with extensive experience in customer service. A dataset example is shown in Appendix E.
The dataset contains 1,574 total dialogue turns with an average of 5.9 turns per dialogue (ranging from 2 to 24 turns) and 22.0 words per utterance on average. All dialogues have been rigorously anonymized to protect user privacy. Following standard practice, we partitioned the dataset into training set ($80\%$, 214 dialogues with 1,299 conversation turns), validation set ($10\%$, 27 dialogues with 149 turns), and test set ($10\%$, 27 dialogues with 126 turns).
@@ 4.1.2 Evaluation Metrics
Automatic Metrics: We calculated BLEU, ROUGE, METEOR and BERTSCORE between system response and ground truth responses, as well as between retrieved responses and ground truth responses.
LLM-as-a-Judge Evaluation: We utilized Claude 3.7 Sonnet [16] as an independent evaluator to assess both retrieved response quality and generated responses. For each evaluation instance, the LLM was provided with the complete dialogue history, the current user query, anonymized system responses generated by different systems, and the ground truth response for reference. For each comparison, LLM compared outputs from all systems across five dimensions (relevance, helpfulness, style consistency, contextual appropriateness, and professionalism) and provided numerical ratings (1-10) for each system, with the highest-rated system designated as "winner" for that example. When multiple systems receive identical highest ratings, all are counted as winners. We employ both average ratings and win counts to analyze system performance. For retrieved quality evaluation, we excluded the "professionalism" dimension, and removed ground truth responses in the LLM prompt. The complete prompt template for response generation is provided in Appendix D.
@@ 4.2 Hyperparameter Analysis
Our empirical analysis reveals two key findings regarding the parameter $\alpha$: (1) the configuration with $\alpha=0.1$ consistently achieves optimal performance across all evaluation metrics, with performance degradation as intent weight increases; and (2) the configurations with $\alpha\geq 0.5$ yield identical retrieval results, suggesting a saturation threshold beyond which intent-based retrieval dominates the ranking process. To validate these automatic evaluation results, we performed LLM-as-a-Judge evaluation using the same criteria from Section 4.1.2. Among the cases where retrieval results differed, $\alpha=0.1$ configuration achieved superior performance in 82.9% cases. This alignment between automatic metrics and LLM-as-a-Judge evaluation validates our parameter selection approach.
@@ 4.3 Evaluation Results
Table 2: Performance comparison across all metrics. Bold values indicates best performance for each metric. The percentage values in parentheses show CID-GraphRAG’s relative improvement over the second-best system.
Retrieval Quality Response Generation
@@
Table 2 presents comprehensive results comparing CID-GraphRAG with baselines for retrieval quality and response generation respectively. To better illustrate the overall performance of different systems in our comparative evaluation, Figure 3 presents the win counts from LLM-as-a-Judge evaluations for both retrieval quality and response generation quality.
Retrieval Quality: CID-GraphRAG demonstrates consistent superiority in retrieval quality across all metrics. For automatic metrics, it achieves relative gains of 2.9% in BLEU-4, 5.2% in ROUGE-L, and a substantial 7.0% in METEOR compared to the second-best system. The LLM-as-a-Judge evaluation reveals even more significant improvements, with CID-GraphRAG outperforming other approaches across all dimensions, particularly in relevance (+12.5%) and helpfulness (+9.3%), demonstrating that our dual-pathway approach retrieves more useful examples. These improvements are further validated by win count analysis (Figure 3), where CID-GraphRAG (73 wins) outperforms the second-best system Conversation RAG (51 wins) by 43%.
@@
LLM-as-a-Judge evaluation of generated responses reveals significant advantages. CID-GraphRAG achieves the highest scores across all dimensions, with notable improvements in style consistency (+4.9%) and professionalism (+4.3%). Most importantly, CID-GraphRAG records the highest number of wins (60 wins) in head-to-head comparisons, representing a 57.9% improvement over the second best (Conversation RAG, 38 wins).
@@
Figure 3: Performance comparison of different methods based on LLM-as-a-Judge win counts. The left group shows retrieval quality wins, and the right group shows response generation wins. CID-GraphRAG consistently outperforms all baseline methods in both aspects, with particularly significant advantages in retrieval quality.
@@ 4.4 Discussion
Comparing human and LLM-as-a-Judge evaluations reveals interesting differences in system rankings. While both confirm CID-GraphRAG’s superiority, LLM-as-a-Judge identified Direct LLM as competitive in style consistency (7.83) and relevance (7.81), whereas human evaluators consistently ranked Intent RAG second, particularly for style consistency (9.38) and professionalism (9.03). This suggests domain experts place higher value on intent-guided responses than LLM evaluators, especially for maintaining professional communication standards in customer service contexts.
Approach Input Tokens Latency (ms)
Intent Generation Intent Retrieval &
Recognition Recognition Generation
CID-GraphRAG 2136 1694 2649 2501
CID-GraphRAG (optimized) - 1694 11 2501
Intent RAG 2136 1542 2649 2467
Conversation RAG - 1702 - 2529
Direct LLM - 1203 - 2320
While CID-GraphRAG delivers superior retrieval and response quality, it introduces additional computational overhead primarily from LLM-based intent recognition, resulting in higher token consumption and latency compared to baseline approaches. To quantify these computational trade-offs, Table 3 presents a detailed analysis of token consumption and latency across all evaluated approaches. This analysis reveals that intent recognition contributes approximately 2,136 tokens and 2.6 seconds to the processing pipeline. To explore potential efficiency improvements, we conducted preliminary timing experiments with a BERT-base classifier for intent recognition. This lightweight approach offers dual advantages: (1) it completely eliminates the 2,136 input tokens required for intent recognition, reducing per-request cost by approximately 56%, and (2) it reduces intent recognition latency from 2.6 seconds to approximately 0.01 seconds (99.6% reduction), bringing CID-GraphRAG’s total latency closer to the Direct LLM baseline. However, the impact of this optimization on retrieval and response quality requires further investigation.
This computational flexibility stems from our framework’s modular architecture, which decouples intent recognition from the dual-retrieval mechanism, allowing for domain-specific intent recognition approaches including fine-tuned models, LLM-based classification, or hybrid solutions, depending on specific business requirements and computational constraints.